# 3cm3cm3cm3cm

### v1

@@ Abstract.
We study the joint action of the Bethe algebra $\mathcal{B}_{d}^{\bf w}$ for the general linear Lie algebra ${\mathfrak{gl}}_{d}$ with respect to ${\bf w}\in{\mathbb{C}}^{d}$ and the Bethe algebra ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ for the general linear Lie superalgebra ${\mathfrak{gl}}_{p+m|q+n}$ with respect to ${\bf z}\in{\mathbb{C}}^{p+q+m+n}$ on the Fock space of $d(p+m)$ bosonic and $d(q+n)$ fermionic oscillators. We establish a duality, called the Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$, which is an equivalence between the actions of the algebras $\mathcal{B}_{d}^{\bf w}$ and ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ on the Fock space. As an application, we show that the action of ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ on each weight space of the evaluation module $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf w})$, where $\underline{M\mkern-4.0mu}\mkern 4.0mu$ is a $d$-fold tensor product of certain classes of infinite-dimensional unitarizable highest weight ${\mathfrak{gl}}_{p+m|q+n}$-modules, is a cyclic ${\mathcal{B}_{p+m|q+n}^{\bf z}}$-module, and that ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ is diagonalizable with a simple spectrum on the weight space for generic ${\bf w}$ and ${\bf z}$.
@@ 1. Introduction
Let $\mathcal{F}:={\mathbb{C}}[{\bf x},{\bf y},{\boldsymbol{\eta}},{\boldsymbol{\zeta}}]$ be the polynomial superalgebra generated by $x^{a}_{i}$, $\eta^{a}_{j}$, $y^{a}_{r}$ and $\zeta^{a}_{s}$, for $i=1,\ldots,m$, $j=1,\ldots,n$, $r=1,\ldots,p$, $s=1,\ldots,q$ and $a=1,\ldots,d$. Here $x^{a}_{i}$ and $y^{a}_{r}$ are even variables while $\eta^{a}_{j}$ and $\zeta^{a}_{s}$ are odd. The superalgebra $\mathcal{F}$ can be realized as the Fock space of $d(p+m)$ bosonic and $d(q+n)$ fermionic oscillators. Let $\mathcal{D}:={\mathbb{D}}[{\bf x},{\bf y},{\boldsymbol{\eta}},{\boldsymbol{\zeta}}]$ be the corresponding Weyl superalgebra, i.e., $\mathcal{D}$ is the superalgebra generated by $x^{a}_{i}$, $\eta^{a}_{j}$, $y^{a}_{r}$ and $\zeta^{a}_{s}$ as well as their derivatives $\frac{\partial}{\partial x_{i}^{a}}$, $\frac{\partial}{\partial\eta_{j}^{a}}$, $\frac{\partial}{\partial y_{r}^{a}}$ and $\frac{\partial}{\partial\zeta_{r}^{a}}$. There is a natural action of $\mathcal{D}$ on $\mathcal{F}$, and hence $\mathcal{D}$ is a subalgebra of the endomorphism algebra ${\rm End}(\mathcal{F})$ of $\mathcal{F}$.
Moreover, there exist superalgebra homomorphisms $\phi:U({\mathfrak{gl}}_{d})\longrightarrow\mathcal{D}$ and $\overline{\phi}:U({\mathfrak{gl}}_{p+m|q+n})\longrightarrow\mathcal{D}$ from the universal enveloping algebras of ${\mathfrak{gl}}_{d}$ and ${\mathfrak{gl}}_{p+m|q+n}$ to $\mathcal{D}$ defined by (4.2) and (4.3), respectively. These maps induce a ${\mathfrak{gl}}_{d}\times{\mathfrak{gl}}_{p+m|q+n}$-action on $\mathcal{F}$. It is well-known that $\mathcal{F}$ decomposes into a direct sum of unitarizable ${\mathfrak{gl}}_{d}\times{\mathfrak{gl}}_{p+m|q+n}$-modules. This is called the Howe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$ ([CLZ]). Such unitarizable modules are, in general, infinite-dimensional. Taking $p=q=0$, the maps $\phi$ and $\overline{\phi}$ specialize to the superalgebra homomorphisms $U({\mathfrak{gl}}_{d})\longrightarrow{\mathbb{D}}[{\bf x},{\boldsymbol{\eta}}]$ and $U({\mathfrak{gl}}_{m|n})\longrightarrow{\mathbb{D}}[{\bf x},{\boldsymbol{\eta}}]$, respectively, and the duality reduces to the Howe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{m|n})$, which yields a decomposition of ${\mathbb{C}}[{\bf x},{\boldsymbol{\eta}}]$ into a direct sum of finite-dimensional ${\mathfrak{gl}}_{d}\times{\mathfrak{gl}}_{m|n}$-modules ([CW, Se1, Se2]). Assuming further that $n=0$, it gives the Howe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{m})$ ([H]).
Fix ${\bf w}\in{\mathbb{C}}^{\ell}$ and ${\bf z}\in{\mathbb{C}}^{p+q+m+n}$. There are natural extensions of $\phi$ and $\overline{\phi}$ to the universal enveloping algebras of current algebras, and we have superalgebra homomorphisms $\phi_{\bf z}:U({\mathfrak{gl}}_{d}[t])\longrightarrow\mathcal{D}$ and $\overline{\phi}_{\bf w}:U({\mathfrak{gl}}_{p+m|q+n}[t])\longrightarrow\mathcal{D}$; see (4.4) and (4.5). It is natural to ask if dualities exist in the context of current algebras. When $p=q=n=0$, a duality phenomenon is noticed in [TL] (see also [TV]). In [MTV3], a precise statement of the duality between $U({\mathfrak{gl}}_{d}[t])$ and $U({\mathfrak{gl}}_{m}[t])$ is made and proved. So far, various generalizations have emerged (see [HM, TU, VY]). All of these are related to the Gaudin models for general linear Lie (super)algebras.
We will establish a duality between $U({\mathfrak{gl}}_{d}[t])$ and $U({\mathfrak{gl}}_{p+m|q+n}[t])$ in this paper. We consider the Bethe algebra $\mathcal{B}_{d}^{\bf w}$ for ${\mathfrak{gl}}_{d}$ with respect to ${\bf w}$ and the Bethe algebra ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ for ${\mathfrak{gl}}_{p+m|q+n}$ with respect to ${\bf z}$; see Section 3.3 and Section 3.4. As the algebras $\mathcal{B}_{d}^{\bf w}$ and ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ are respectively subalgebras of $U({\mathfrak{gl}}_{d}[t])$ and $U({\mathfrak{gl}}_{p+m|q+n}[t])$, the maps $\phi_{\bf z}$ and $\overline{\phi}_{\bf w}$ induce an action of $\mathcal{B}_{d}^{\bf w}$ and an action of ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ on the polynomial superalgebra $\mathcal{F}$, respectively. We show that these two actions are equivalent. More precisely, we have:
@@
$\phi_{\bf z}\big{(}\mathcal{B}_{d}^{\bf w}\big{)}=\overline{\phi}_{\bf w}\big{(}{\mathcal{B}_{p+m|q+n}^{\bf z}}\big{)}$.
@@
We name Theorem 1.1 the Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$. It recovers several dualities established earlier. After specializing to $p=q=0$, it recovers the Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{m|n})$ due to Huang and Mukhin [HM]. If we assume further that $n=0$ (resp., $m=0$), Theorem 1.1 gives the Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{m})$ obtained by Mukhin, Tarasov, and Varchenko [MTV3] (resp., a variant of the duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{n})$ obtained by Tarasov and Uvarov [TU]). Furthermore, we will show in [ChL3] that there is another duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$ for Gaudin models with irregular singularities, which specializes to a variant of Theorem 1.1.
@@
The Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$ enables us to study the action of ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ on $\mathcal{F}$ via the action of $\mathcal{B}_{d}^{\bf w}$ on $\mathcal{F}$, and vice versa. Using the properties of the action of $\mathcal{B}_{d}^{\bf w}$ on finite-dimensional ${\mathfrak{gl}}_{d}$-modules, we will better understand the action of ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ on unitarizable ${\mathfrak{gl}}_{p+m|q+n}$-modules. Let $\underline{M\mkern-4.0mu}\mkern 4.0mu:=M_{1}\otimes\cdots\otimes M_{d}$, where each $M_{i}$ is an infinite-dimensional unitarizable highest weight ${\mathfrak{gl}}_{p+m|q+n}$-module with highest weight corresponding to a generalized partition of depth 1 (see Section 5.2 for a detailed description). We will see in Theorem 5.1 that each weight space of the evaluation ${\mathfrak{gl}}_{p+m|q+n}[t]$-module $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf w})$ is a cyclic ${\mathcal{B}_{p+m|q+n}^{\bf z}}$-module, and that ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ is diagonalizable with a simple spectrum on the weight space for generic ${\bf w}$ and ${\bf z}$.
Let us explain a little how we prove Theorem 1.1. The Bethe algebras $\mathcal{B}_{d}^{\bf w}$ and ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ are determined by the column determinant ${\rm cdet}\,({\mathcal{L}}_{d}^{\bf w})$ and the Berezinian ${\rm Ber}^{\bf s}\,({\mathcal{L}}_{{p+m|q+n}}^{\bf z})$ of type ${\bf s}:=(0^{p},1^{q},0^{m},1^{n})$, respectively; see (3.6) and (3.9). In the special case where $p=q=0$, the specializations of the series $\phi_{\bf z}({\rm cdet}\,({\mathcal{L}}_{d}^{\bf w}))$ and $\overline{\phi}_{\bf w}({\rm Ber}^{\bf s}\,({\mathcal{L}}_{{p+m|q+n}}^{\bf z}))$ admit Capelli-type identities [HM, Proposition 5.4 and Proposition 5.6], and the duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{m|n})$ is obtained by matching the identities. However, the situation is somewhat complicated in our case as no such identities are known to exist. Based on the properties of Berezinians, we can relate $\overline{\phi}_{\bf w}({\rm Ber}^{\bf s}\,({\mathcal{L}}_{{p+m|q+n}}^{\bf z}))$ to $\phi_{\bf z}({\rm cdet}\,({\mathcal{L}}_{d}^{\bf w}))$ (for general $p,q,m,n$) by examining an anti-involution on the Weyl superalgebra $\mathcal{D}$ and an anti-involution on the Bethe algebra $\mathcal{B}_{d}^{\bf w}$ (see Proposition 4.5). Theorem 1.1 will then follow.
We organize the paper as follows. In Section 2, we set up our notations. We recall the notions of the universal affine vertex algebra at the critical level, the Feigin–Frenkel center and the Bethe algebra. In Section 3, we discuss the Bethe algebras for general linear Lie (super)algebras and their basic properties. In Section 4, we introduce the joint action of the Bethe algebras $\mathcal{B}_{d}^{\bf w}$ and ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ on the polynomial superalgebra $\mathcal{F}$ for any sequences ${\bf w}\in{\mathbb{C}}^{\ell}$ and ${\bf z}\in{\mathbb{C}}^{p+q+m+n}$. We prove Theorem 1.1, i.e., the Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$. In Section 5, we apply the Bethe duality to the action of ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ on unitarizable ${\mathfrak{gl}}_{p+m|q+n}$-modules and obtain Theorem 5.1.

### v2

# Bethe algebras for unitarizable modules over classical Lie (super)algebras and a duality

@@ Abstract.
Let $\mathfrak{g}$} denote the classical Lie algebra ${\mathfrak{gl}}_{d}$, $\mathfrak{sp}_{2d}$, or $\mathfrak{so}_{2d}$ with a fixed *-structure $\bm{\sigma}$}. Let $M_{1},\ldots,M_{\ell}$ be unitarizable $\mathfrak{g}$}-modules (with respect to $\bm{\sigma}$}), and let ${\bf z}=(z_{1},\ldots,z_{\ell})\in{\mathbb{C}}^{\ell}$. We investigate the action of the Bethe algebra $\mathcal{B}_{\mathfrak{g}}^{\mu}$ for $\mathfrak{g}$} with respect to $\mu\in{\mathfrak{g}}^{*}$ on the tensor product $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf z}):=M_{1}(z_{1})\otimes\cdots\otimes M_{\ell}(z_{\ell})$ of evaluation ${\mathfrak{g}}[t]$-modules. We show that if $\mu\circ{\bm{\sigma}}$ equals the complex conjugation of $\mu$, then $\mathcal{B}_{\mathfrak{g}}^{\mu}$ is diagonalizable on any finite-dimensional $\mathcal{B}_{\mathfrak{g}}^{\mu}$-submodule of $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf z})$ for $\bf z$$\in{\mathbb{R}}^{\ell}$. This, together with the result derived from the duality of Bethe algebras (see below), suggests that a simple spectrum conjecture for $\mathcal{B}_{\mathfrak{g}}^{\mu}$ should hold.
We establish a duality of Bethe algebras for the general linear Lie (super)algebras ${\mathfrak{gl}}_{d}$ and ${\mathfrak{gl}}_{p+m|q+n}$. As an application, we show that under a generic condition, the Bethe algebra for ${\mathfrak{gl}}_{p+m|q+n}$ with respect to ${\bf z}\in{\mathbb{C}}^{p+q+m+n}$ is diagonalizable with a simple spectrum on any weight space of $L_{1}(w_{1})\otimes\cdots\otimes L_{d}(w_{d})$, where the $L_{i}$ are (infinite-dimensional) unitarizable highest weight ${\mathfrak{gl}}_{p+m|q+n}$-modules corresponding to generalized partitions of depth 1, and $w_{1},\ldots,w_{d}\in{\mathbb{C}}$. We also obtain the corresponding result for ${\mathfrak{gl}}_{p+m}$ by setting $q=n=0$.
@@ 1. Introduction
Let $\mathfrak{g}$} be a finite-dimensional semisimple Lie algebra over $\mathbb{C}$}. The Bethe algebra for $\mathfrak{g}$} is a commutative subalgebra of $U({\mathfrak{g}}[t])$ induced by the Feigin–Frenkel center (see [FFR, FFTL, Ry1]), where $U({\mathfrak{g}}[t])$ denotes the universal enveloping algebra of the current algebra ${\mathfrak{g}}[t]$. It produces (higher) Gaudin Hamiltonians, including the quadratic ones originally introduced by Gaudin [G1, G2] to describe a completely integrable quantum spin chain model. Also, the notion of the Bethe algebra can be extended to the general linear Lie (super)algebras (see [HM, MR, MTV1, Ta]).
Let $d$ be a positive integer, and let $p$, $q$, $m$, and $n$ be non-negative integers that are not all zero. The aim of the paper is two-fold. Firstly, we investigate the action of the Bethe algebra for $\mathfrak{g}$} on unitarizable $\mathfrak{g}$}-modules, where $\mathfrak{g}$} is the classical Lie algebra of type $\mathfrak{a,c,d}$ of rank $d$ (i.e., ${\mathfrak{g}}={\mathfrak{gl}}_{d}$, $\mathfrak{sp}_{2d}$, or $\mathfrak{so}_{2d}$). Secondly, we establish a duality of Bethe algebras for the general linear Lie (super)algebras ${\mathfrak{gl}}_{d}$ and ${\mathfrak{gl}}_{p+m|q+n}$ and use it to study the action of the Bethe algebra for ${\mathfrak{gl}}_{p+m|q+n}$ on some infinite-dimensional unitarizable ${\mathfrak{gl}}_{p+m|q+n}$-modules.
1.1. Unitarizable modules and diagonalization of Bethe algebras
Let $\mathfrak{g}$} be a simple Lie algebra or the general linear Lie algebra over $\mathbb{C}$}. For $\mu\in{\mathfrak{g}}^{*}$, let $\mathcal{B}_{\mathfrak{g}}^{\mu}$ be the Bethe algebra for $\mathfrak{g}$} with respect to $\mu$; see Section 3.2. The action of $\mathcal{B}_{\mathfrak{g}}^{\mu}$ on a tensor product of finite-dimensional $\mathfrak{g}$}-modules has been extensively studied (see [FFRy, MTV1, MTV2, MTV4, MTV5, Ry3]). Let ${\bf z}=(z_{1},\ldots,z_{\ell})\in{\mathbb{C}}^{\ell}$, and let $V_{1},\ldots,V_{\ell}$ be finite-dimensional irreducible $\mathfrak{g}$}-modules. Feigin, Frenkel, and Rybnikov [FFRy] show that the algebra $\mathcal{B}_{\mathfrak{g}}^{\mu}$ acts cyclically on the tensor product $\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z}):=V_{1}(z_{1})\otimes\cdots\otimes V_{\ell}(z_{\ell})$ of evaluation ${\mathfrak{g}}[t]$-modules for any regular $\mu$ and any pairwise distinct $z_{1},\ldots,z_{\ell}$. They also show that $\mathcal{B}_{\mathfrak{g}}^{\mu}$ is diagonalizable with a simple spectrum on $\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})$ for generic $\mu$ and generic $\bf z$}.
Now let ${\mathfrak{g}}={\mathfrak{gl}}_{d}$, $\mathfrak{sp}_{2d}$, or $\mathfrak{so}_{2d}$. It possesses a *-structure ${\bm{\sigma}}:{\mathfrak{g}}\longrightarrow{\mathfrak{g}}$, i.e., $\bm{\sigma}$} is an even anti-linear anti-involution (see Section 3.3). Let $M_{1},\ldots,M_{\ell}$ be unitarizable $\mathfrak{g}$}-modules with respect to $\bm{\sigma}$}, and let $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf z})=M_{1}(z_{1})\otimes\cdots\otimes M_{\ell}(z_{\ell})$. Define ${\mathfrak{g}}^{*}_{{\bm{\sigma}}}=\left\{\,\mu\in{\mathfrak{g}}^{*}\;\middle|\;\mu\circ{\bm{\sigma}}=\overline{\mu}\,\right\}$, where $\overline{\mu}\in{\mathfrak{g}}^{*}$ is the complex conjugation of $\mu$. Using the properties of *-structures and Feigin–Frenkel centers, we obtain the following.
@@ 1.1. Unitarizable modules and diagonalization of Bethe algebras
The Bethe algebra $\mathcal{B}_{\mathfrak{g}}^{\mu}$ is diagonalizable on any finite-dimensional $\mathcal{B}_{\mathfrak{g}}^{\mu}$-submodule of $\underline{M\mkern-4.0mu}\mkern 4.0mu({\bf z})$ for any $\mu\in{\mathfrak{g}}^{*}_{{\bm{\sigma}}}$ and $\bf z$$\in{\mathbb{R}}^{\ell}$.
For ${\bf w}:=(w_{1},\ldots,w_{d})\in{\mathbb{C}}^{d}$, let $\mu_{\bf w}\in{\mathfrak{g}}^{*}$ be the element defined by (3.12) and (3.13). In other words, $\mu_{\bf w}$ corresponds to a diagonal matrix. We write $\mathcal{B}_{\mathfrak{g}}^{\bf w}:=\mathcal{B}_{\mathfrak{g}}^{\mu_{\bf w}}$.
Let $L_{1},\ldots,L_{\ell}$ be unitarizable highest weight $\mathfrak{g}$}-modules with respect to $\bm{\sigma}$}. Set $\underline{L}=L_{1}\otimes\cdots\otimes L_{\ell}$ and $\underline{L}({\bf z})=L_{1}(z_{1})\otimes\cdots\otimes L_{\ell}(z_{\ell})$. Let $\mathfrak{h}$} denote the standard Cartan subalgebra of $\mathfrak{g}$} (see Section 2.1.4). Note that for any weight $\gamma\in{\mathfrak{h}}^{*}$ of $\underline{L}$, the $\gamma$-weight space $\underline{L}({\bf z})_{\gamma}$ of $\underline{L}({\bf z})$ is a finite-dimensional $\mathcal{B}_{\mathfrak{g}}^{\bf w}$-module, and that $\mu_{\bf w}\in{\mathfrak{g}}^{*}_{{\bm{\sigma}}}$ for ${\bf w}\in{\mathbb{R}}^{d}$. We have the following as a consequence of Theorem 1.1.
Corollary 1.2 (Corollary 3.12).
Let $\gamma\in{\mathfrak{h}}^{*}$ be any weight of $\underline{L}$. Then $\mathcal{B}_{\mathfrak{g}}^{\bf w}$ is diagonalizable on $\underline{L}({\bf z})_{\gamma}$ for any ${\bf w}\in{\mathbb{R}}^{d}$ and $\bf z$$\in{\mathbb{R}}^{\ell}$.
A natural question is whether there are unitarizable analogs of the results of [FFRy] mentioned above.
Conjecture 1.3 (Conjecture 3.13).
Let $\gamma\in{\mathfrak{h}}^{*}$ be any weight of $\underline{L}$. Then the weight space $\underline{L}({\bf z})_{\gamma}$ is a cyclic $\mathcal{B}_{\mathfrak{g}}^{\bf w}$-module for any pairwise distinct $w_{1},\ldots,w_{d}\in{\mathbb{C}}$ and any pairwise distinct $z_{1},\ldots,z_{\ell}\in{\mathbb{C}}$. Moreover, $\mathcal{B}_{\mathfrak{g}}^{\bf w}$ is diagonalizable with a simple spectrum on $\underline{L}({\bf z})_{\gamma}$ for generic $\bf w$} and generic $\bf z$}.
Theorem 1.5 (see below) provides positive evidence for Conjecture 1.3 when $q$ and $n$ are set to 0.
1.2. A duality for the general linear Lie (super)algebras
For ${\bf w}:=(w_{1},\ldots,w_{d})\in{\mathbb{C}}^{d}$ and ${\bf z}:=(z_{1},\ldots,z_{p+q+m+n})\in{\mathbb{C}}^{p+q+m+n}$, we consider the Bethe algebra $\mathcal{B}_{d}^{\bf w}:=\mathcal{B}_{{\mathfrak{gl}}_{d}}^{\bf w}$ for the general linear Lie algebra ${\mathfrak{gl}}_{d}$ with respect to $\bf w$} and the Bethe algebra ${\mathcal{B}_{p+m|q+n}^{\bf z}}:={\mathcal{B}_{{\mathfrak{gl}}_{p+m|q+n}}^{\bf z}}$ for the general linear Lie superalgebra ${\mathfrak{gl}}_{p+m|q+n}$ with respect to $\bf z$}; see Section 4.2. Note that $\mathcal{B}_{d}^{\bf w}$ is a subalgebra of $U({\mathfrak{gl}}_{d}[t])$ generated by the coefficients of the series ${\rm cdet}({\mathcal{L}}_{d}^{\bf w})$ while $\mathcal{B}_{p+m|q+n}^{\bf z}$} is a subalgebra of $U({\mathfrak{gl}}_{p+m|q+n}[t])$ generated by the coefficients of the series ${\rm Ber}^{\bf s}\big({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\big)$. Here $\rm cdet$} and ${\rm Ber}^{\bf s}$ are respectively the column determinant and the Berezinian of type ${\bf s}:=(0^{p},1^{q},0^{m},1^{n})$ (see Section 4.1), and ${\mathcal{L}}_{d}^{\bf w}$ and ${\mathcal{L}}_{{p+m|q+n}}^{\bf z}$ are respectively the matrices defined in (4.4) and (4.7).
Let $\mathcal{F}$ be the polynomial superalgebra generated by $x^{a}_{i}$, $\eta^{a}_{j}$, $y^{a}_{r}$, and $\zeta^{a}_{s}$, for $i=1,\ldots,m$, $j=1,\ldots,n$, $r=1,\ldots,p$, $s=1,\ldots,q$, and $a=1,\ldots,d$. Here $x^{a}_{i}$ and $y^{a}_{r}$ are even variables while $\eta^{a}_{j}$ and $\zeta^{a}_{s}$ are odd. The superalgebra $\mathcal{F}$ can be realized as the Fock space of $d(p+m)$ bosonic and $d(q+n)$ fermionic oscillators (cf. [CL, LZ]). Let $\mathcal{D}$ be the corresponding Weyl superalgebra, i.e., $\mathcal{D}$ is the associative unital superalgebra generated by $x^{a}_{i}$, $\eta^{a}_{j}$, $y^{a}_{r}$, and $\zeta^{a}_{s}$ as well as their derivatives $\frac{\partial}{\partial x_{i}^{a}}$, $\frac{\partial}{\partial\eta_{j}^{a}}$, $\frac{\partial}{\partial y_{r}^{a}}$, and $\frac{\partial}{\partial\zeta_{r}^{a}}$. There exist superalgebra homomorphisms $\phi:U({\mathfrak{gl}}_{d})\longrightarrow\mathcal{D}$ and $\varphi:U({\mathfrak{gl}}_{p+m|q+n})\longrightarrow\mathcal{D}$; see (4.9) and (4.10). They extend to superalgebra homomorphisms $\phi_{\bf z}:U({\mathfrak{gl}}_{d}[t])(\!(z^{-1},\partial_{z}^{-1})\!)\longrightarrow\mathcal{D}(\!(z^{-1},\partial_{z}^{-1})\!)$ and $\varphi_{\bf w}:U({\mathfrak{gl}}_{p+m|q+n}[t])(\!(z^{-1},\partial_{z}^{-1})\!)\longrightarrow\mathcal{D}(\!(z^{-1},\partial_{z}^{-1})\!)$, respectively. Here, for instance, $\mathcal{D}(\!(z^{-1},\partial_{z}^{-1})\!)$ is the superalgebra of pseudo-differential operators over $\mathcal{D}$; see Section 2.3. The maps $\phi_{\bf z}$ and $\varphi_{\bf w}$ induce an action of $\mathcal{B}_{d}^{\bf w}$ and an action of $\mathcal{B}_{p+m|q+n}^{\bf z}$} on $\mathcal{F}$, respectively. The following theorem asserts that the actions are equivalent. We call it the Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$.
Theorem 1.4 (Theorem 4.10).
The coefficients of $\phi_{\bf z}{\rm cdet}({\mathcal{L}}_{d}^{\bf w})$ are determined by those of $\varphi_{\bf w}{\rm Ber}^{\bf s}\big({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\big)$, and vice versa. Consequently, $\phi_{\bf z}\big(\mathcal{B}_{d}^{\bf w}\big)=\varphi_{\bf w}\big({\mathcal{B}_{p+m|q+n}^{\bf z}}\big)$.
@@ 1.2. A duality for the general linear Lie (super)algebras
Theorem 1.4 recovers several dualities established earlier. After specializing to $p=q=0$, it recovers the Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{m|n})$ due to Huang and Mukhin [HM]. If we assume further that $n=0$ (resp., $m=0$), Theorem 4.10 gives the Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{m})$ obtained by Mukhin, Tarasov, and Varchenko [MTV3] (resp., a variant of the duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{n})$ obtained by Tarasov and Uvarov [TU]).
Also, another duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{m})$ is obtained by Vicedo and Young [VY] for Gaudin models with irregular singularities. It is extended by the authors [ChL3] to a duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$, which specializes to a variant of Theorem 1.4.
Using the properties of the actions of $\mathcal{B}_{d}^{\bf w}$ on tensor products of finite-dimensional ${\mathfrak{gl}}_{d}$-modules, Theorem 1.4 enables us to better understand the actions of $\mathcal{B}_{p+m|q+n}^{\bf z}$} on tensor products of infinite-dimensional unitarizable ${\mathfrak{gl}}_{p+m|q+n}$-modules, as shown in the following theorem.
Theorem 1.5 (Theorem 4.13).
@@
Let $\underline{L}=L_{1}\otimes\cdots\otimes L_{d}$ and $\underline{L}({\bf w})=L_{1}(w_{1})\otimes\cdots\otimes L_{d}(w_{d})$, where the $L_{i}$ are (infinite-dimensional) unitarizable highest weight ${\mathfrak{gl}}_{p+m|q+n}$-modules corresponding to generalized partitions of depth 1 (see (4.17)). Let $\mu\in{\mathfrak{h}}_{{p+m|q+n}}^{*}$ be any weight of $\underline{L}$ (where ${\mathfrak{h}}_{{p+m|q+n}}$ is the Cartan subalgebra of ${\mathfrak{gl}}_{p+m|q+n}$ defined in Section 2.1.1). Then the weight space $\underline{L}({\bf w})_{\mu}$ is a cyclic $\mathcal{B}_{p+m|q+n}^{\bf z}$}-module for any pairwise distinct $z_{1},\ldots,z_{p+q+m+n}\in{\mathbb{C}}$ and any pairwise distinct $w_{1},\ldots,w_{d}\in{\mathbb{C}}$. Moreover, $\mathcal{B}_{p+m|q+n}^{\bf z}$} is diagonalizable with a simple spectrum on $\underline{L}({\bf w})_{\mu}$ for generic $\bf z$} and generic $\bf w$}.
1.3. Outline
We organize the paper as follows. Section 2 is devoted to reviewing the classical Lie (super)algebras of types $\mathfrak{a,c,d}$ and some background material needed in this paper. In Section 3, we investigate the action of the Bethe algebra for $\mathfrak{g}$} on a tensor product of evaluation ${\mathfrak{g}}[t]$-modules, where $\mathfrak{g}$} is the classical Lie algebra of type $\mathfrak{a,c,d}$. We prove Theorem 1.1 and obtain Corollary 1.2. In Section 4, we establish Theorem 1.4, i.e., the Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$. We give an application of the duality and demonstrate Theorem 1.5, which provides positive evidence for Conjecture 1.3. In Appendix A, we prove Proposition 4.7, which gives an equality relating $\phi_{\bf z}{\rm cdet}({\mathcal{L}}_{d}^{\bf w})$ to $\varphi_{\bf w}{\rm Ber}^{\bf s}\big({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\big)$.